# Conformal Prediction for Ensembles: Improving Efficiency via Score-Based Aggregation

### v2

@@ Appendix D Additional UCI Results
Table 5: Average coverages across tasks for $\alpha=0.05$ are shown in the top row and average prediction set lengths in the bottom row, where both were assessed over a batch of i.i.d. test samples (10% of the dataset size). Standard deviations and means were computed across 10 randomizations of draws of the training, calibration, and test sets. Note that, while the single-stage prediction regions are the smallest, they fail to achieve the desired coverage level and are, therefore, precluded from comparison.
@@
Dataset Methods Linear Model LASSO Random Forest Neural Net $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA (Single-Stage) CSA
3droad Coverage 0.947 (0.001) 0.947 (0.001) 0.948 (0.001) 0.95 (0.0) 0.945 (0.001) 0.915 (0.0) 0.948 (0.0) 0.926 (0.001) 0.95 (0.0)
Length 71.583 (0.325) 71.576 (0.331) 5.362 (0.006) 56.298 (0.602) 53.166 (0.482) 29.04 (0.245) 51.158 (0.015) 1.175 (0.036) 1.361 (0.031)
airfoil Coverage 0.95 (0.012) 0.981 (0.018) 0.954 (0.02) 0.969 (0.014) 0.962 (0.014) 0.939 (0.024) 0.969 (0.014) 0.854 (0.04) 0.952 (0.006)
Length 18.742 (0.313) 25.218 (1.05) 7.353 (0.672) 25.256 (1.424) 18.097 (0.565) 11.96 (0.634) 18.303 (0.787) 1.873 (0.21) 2.714 (0.141)
bike Coverage 0.963 (0.004) 0.959 (0.004) 0.948 (0.001) 0.955 (0.003) 0.955 (0.003) 0.932 (0.008) 0.959 (0.003) 0.946 (0.005) 0.955 (0.003)
Length 3.397 (0.131) 3.416 (0.123) 0.025 (0.0) 1.716 (0.104) 1.654 (0.11) 0.837 (0.055) 2.094 (0.09) 0.569 (0.097) 0.629 (0.105)
parkinsons Coverage 0.948 (0.002) 0.951 (0.001) 0.935 (0.005) 0.942 (0.001) 0.946 (0.002) 0.878 (0.006) 0.937 (0.005) 0.857 (0.019) 0.961 (0.004)
Length 37.058 (0.599) 37.349 (0.34) 0.876 (0.11) 18.912 (0.931) 17.918 (0.814) 8.738 (0.427) 23.945 (0.039) 0.465 (0.196) 0.588 (0.188)
pol Coverage 0.942 (0.007) 0.944 (0.007) 0.951 (0.002) 0.952 (0.001) 0.936 (0.006) 0.9 (0.009) 0.949 (0.003) 0.889 (0.007) 0.95 (0.003)
Length 97.558 (0.808) 97.476 (1.0) 19.437 (0.374) 41.775 (1.306) 38.033 (1.092) 27.596 (0.77) 63.011 (0.894) 1.357 (0.019) 1.654 (0.032)
pumadyn32nm Coverage 0.932 (0.008) 0.933 (0.008) 0.943 (0.004) 0.924 (0.011) 0.917 (0.008) 0.886 (0.013) 0.935 (0.008) 0.863 (0.018) 0.926 (0.006)
Length 3.865 (0.043) 3.827 (0.061) 1.158 (0.005) 1.145 (0.012) 1.239 (0.011) 1.089 (0.013) 2.54 (0.029) 2.355 (0.124) 2.767 (0.076)
slice Coverage 0.951 (0.001) 0.948 (0.001) 0.948 (0.0) 0.953 (0.003) 0.947 (0.0) 0.912 (0.0) 0.951 (0.002) 0.91 (0.004) 0.95 (0.001)
Length 35.242 (0.099) 35.863 (0.057) 2.749 (0.059) 8.596 (0.412) 8.315 (0.382) 5.378 (0.216) 20.401 (0.162) 1.114 (0.025) 1.388 (0.02)
Appendix E Robust Traffic Routing Setup
@@
Appendix F Conformal Aggregation Methods
We now describe the methods from [25] that were compared against experimentally, specifically the standard majority-vote $\mathcal{C}^{M}$, partially randomized thresholding $\mathcal{C}^{R}$, and fully randomized thresholding $\mathcal{C}^{U}$ approaches. As discussed in Section 2.4, these methods all follow the structural form of Equation 2 and largely differ in their choice of weights and thresholds. The standard majority-vote $\mathcal{C}^{M}$ is the most natural choice, defined by

### v3

@@ Appendix F Full ImageNet Results
Table 5: Average coverages across different coverage levels are shown in the top rows and average prediction set sizes in the bottom rows. Both were assessed over a batch of i.i.d. test samples (15% of the validation set from ImageNet). Standard deviations and means were computed across 10 randomized draws of the calibration and test sets.
Dataset/$\alpha$ Metrics ResNet VGG DenseNet VFCP $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ Ensemble CSA (Single-Stage) CSA
ImageNet Coverage 0.97 (0.011) 0.966 (0.011) 0.939 (0.002) 0.954 (0.006) 0.956 (0.011) 0.948 (0.01) 0.96 (0.013) 0.95 (0.006) 0.955 (0.013) 0.957 (0.01)
$(\alpha=0.07)$ Length 174.828 (2.037) 181.58 (2.459) 160.623 (2.828) 61.247 (1.638) 122.899 (2.154) 111.602 (2.155) 173.264 (2.674) 86.955 (1.886) 44.787 (1.198) 45.352 (1.56)
Coverage 0.95 (0.003) 0.949 (0.004) 0.952 (0.002) 0.95 (0.003) 0.975 (0.002) 0.954 (0.004) 0.95 (0.003) 0.949 (0.002) 0.95 (0.002) 0.95 (0.003)
$(\alpha=0.05)$ Length 220.022 (2.072) 229.523 (3.076) 208.658 (2.016) 78.108 (2.004) 166.933 (2.157) 143.323 (2.932) 220.491 (2.773) 112.161 (2.115) 58.424 (1.674) 59.574 (3.382)
Coverage 0.98 (0.002) 0.981 (0.002) 0.98 (0.002) 0.98 (0.002) 0.992 (0.002) 0.982 (0.002) 0.98 (0.002) 0.98 (0.003) 0.98 (0.001) 0.979 (0.002)
$(\alpha=0.02)$ Length 357.487 (5.046) 363.376 (4.916) 342.479 (5.603) 137.356 (3.595) 311.521 (4.053) 327.754 (6.251) 355.701 (4.423) 202.573 (2.948) 117.272 (5.015) 121.477 (19.719)
Coverage 0.99 (0.001) 0.991 (0.001) 0.989 (0.002) 0.99 (0.001) 0.997 (0.001) 0.991 (0.002) 0.99 (0.002) 0.99 (0.002) 0.99 (0.001) 0.99 (0.002)
$(\alpha=0.01)$ Length 491.952 (6.353) 726.028 (12.157) 459.399 (6.739) 194.691 (4.579) 580.592 (7.715) 532.155 (24.829) 559.188 (7.07) 299.453 (6.526) 180.534 (8.468) 201.32 (46.509)
Appendix G Full OpenML Results
Table 6: Average coverages across tasks for $\alpha=0.05$ are shown in the top row and average prediction set lengths in the bottom row, where both were assessed over a batch of i.i.d. test samples (20% of the dataset size). Standard deviations and means were computed across 5 randomizations of draws of the training, calibration, and test sets. In cases where the method failed to achieve sufficient coverage (defined as $<0.93$), we do not include it in comparison for set length. Similarly, the single-stage approach fails to achieve coverage due to lack of exchangeability with test points.
@@ Appendix G Full OpenML Results
Dataset Metrics Linear Model LASSO Random Forest XGBoost $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ Ensemble CSA (Single-Stage) CSA
361234 Coverage 0.97 (0.011) 0.966 (0.011) 0.939 (0.002) 0.954 (0.006) 0.956 (0.011) 0.948 (0.01) 0.96 (0.013) 0.95 (0.006) 0.955 (0.013) 0.957 (0.01)
Length 9.673 (0.160) 9.645 (0.154) 10.080 (0.160) 9.157 (0.052) 9.196 (0.123) 8.703 (0.086) 9.524 (0.056) 17.759 (0.275) 7.646 (0.073) 7.688 (0.181)
361235 Coverage 0.947 (0.0) 0.945 (0.005) 0.968 (0.016) 0.95 (0.005) 0.955 (0.016) 0.897 (0.005) 0.953 (0.011) 0.932 (0.021) 0.745 (0.011) 0.984 (0.005)
Length 20.961 (0.651) 24.241 (0.246) 10.096 (0.587) 11.387 (0.452) 11.782 (0.057) 16.088 (0.118) 15.823 (1.272) 6.162 (0.458) 11.695 (0.266)
361236 Coverage 0.975 (0.008) 0.975 (0.008) 0.961 (0.0) 0.948 (0.012) 0.948 (0.012) 0.938 (0.012) 0.965 (0.008) 0.934 (0.004) 0.94 (0.004) 0.963 (0.004)
Length 44407.071 (1173.758) 44509.269 (1229.817) 50820.568 (385.951) 41045.069 (1221.808) 43185.942 (1002.516) 40905.295 (1089.411) 44437.938 (851.862) 60509.250 (2410.320) 30953.589 (2482.065) 33439.322 (1275.213)
361237 Coverage 0.969 (0.023) 0.969 (0.023) 0.981 (0.0) 0.923 (0.0) 0.954 (0.015) 0.9 (0.008) 0.969 (0.023) 0.885 (0.038) 0.8 (0.015) 0.977 (0.008)
Length 44.019 (0.990) 44.069 (1.115) 27.035 (1.014) 26.524 (1.244) 31.967 (1.118) 14.473 (0.503) 23.145 (0.199)
361241 Coverage 0.954 (0.001) 0.956 (0.001) 0.944 (0.005) 0.957 (0.002) 0.954 (0.002) 0.923 (0.0) 0.952 (0.0) 0.949 (0.001) 0.917 (0.006) 0.951 (0.001)
Length 19.133 (0.062) 20.245 (0.095) 18.102 (0.055) 18.482 (0.062) 17.958 (0.062) 18.932 (0.034) 29.548 (0.191) 15.199 (0.427) 17.328 (0.097)
361242 Coverage 0.944 (0.004) 0.955 (0.0) 0.947 (0.004) 0.942 (0.0) 0.948 (0.0) 0.914 (0.003) 0.944 (0.001) 0.949 (0.003) 0.9 (0.006) 0.944 (0.002)
Length 70.248 (0.304) 84.510 (0.282) 50.442 (0.421) 54.844 (0.036) 54.217 (0.115) 65.635 (0.094) 61.613 (0.372) 44.602 (0.170) 57.935 (0.070)
361243 Coverage 0.922 (0.03) 0.952 (0.015) 0.956 (0.022) 0.952 (0.015) 0.937 (0.022) 0.893 (0.044) 0.937 (0.022) 0.919 (0.022) 0.748 (0.126) 0.956 (0.022)
Length 71.388 (0.152) 75.924 (2.291) 72.877 (0.729) 68.493 (0.993) 72.048 (0.024) 43.742 (10.285) 68.220 (1.605)
361244 Coverage 0.97 (0.022) 0.97 (0.022) 0.97 (0.022) 0.97 (0.022) 0.97 (0.022) 0.97 (0.022) 0.97 (0.022) 0.974 (0.015) 0.963 (0.037) 0.956 (0.015)
Length 3.274 (0.004) 3.274 (0.004) 3.336 (0.023) 3.284 (0.010) 3.272 (0.000) 3.269 (0.003) 3.289 (0.002) 4.854 (0.293) 0.287 (0.008) 0.287 (0.008)
361247 Coverage 0.96 (0.001) 0.953 (0.003) 0.94 (0.001) 0.951 (0.003) 0.963 (0.001) 0.903 (0.007) 0.951 (0.0) 0.954 (0.001) 0.843 (0.003) 0.943 (0.006)
Length 0.025 (0.000) 0.038 (0.000) 0.006 (0.000) 0.016 (0.000) 0.015 (0.000) 0.022 (0.000) 0.013 (0.000) 0.005 (0.000) 0.008 (0.000)
361249 Coverage 0.96 (0.002) 0.956 (0.002) 0.962 (0.003) 0.972 (0.007) 0.953 (0.005) 0.938 (0.002) 0.965 (0.005) 0.936 (0.01) 0.931 (0.0) 0.953 (0.005)
Length 3.008 (0.006) 3.068 (0.009) 2.800 (0.000) 2.780 (0.025) 2.775 (0.006) 2.558 (0.019) 2.894 (0.000) 4.706 (0.099) 2.216 (0.020) 2.614 (0.043)
Table 7: Average coverages across tasks for $\alpha=0.025$ are shown in the top row and average prediction set lengths in the bottom row, where both were assessed over a batch of i.i.d. test samples (20% of the dataset size). Standard deviations and means were computed across 5 randomizations of draws of the training, calibration, and test sets. In cases where the method failed to achieve sufficient coverage (defined as $<0.96$), we do not include it in comparison for set length. Similarly, the single-stage approach fails to achieve coverage due to lack of exchangeability with test points.
Dataset Metrics Linear Model LASSO Random Forest XGBoost $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ Ensemble CSA (Single-Stage) CSA
361234 Coverage 0.987 (0.008) 0.987 (0.008) 0.974 (0.004) 0.977 (0.008) 0.982 (0.008) 0.971 (0.01) 0.981 (0.01) 0.97 (0.008) 0.976 (0.01) 0.973 (0.006)
Length 11.939 (0.137) 11.871 (0.084) 12.484 (0.168) 11.972 (0.009) 11.587 (0.110) 11.157 (0.086) 11.965 (0.050) 25.598 (0.974) 9.306 (0.259) 8.855 (0.059)
361235 Coverage 0.987 (0.0) 0.982 (0.011) 0.979 (0.011) 0.984 (0.005) 0.989 (0.005) 0.966 (0.016) 0.976 (0.005) 0.958 (0.021) 0.889 (0.011) 0.989 (0.005)
Length 24.595 (0.825) 28.841 (1.129) 11.811 (0.992) 14.237 (0.786) 14.472 (0.172) 12.278 (0.026) 19.231 (0.356) 7.719 (0.467) 12.563 (0.766)
361236 Coverage 0.992 (0.004) 0.992 (0.004) 0.981 (0.0) 0.965 (0.008) 0.975 (0.008) 0.965 (0.008) 0.977 (0.012) 0.955 (0.008) 0.955 (0.012) 0.973 (0.004)
Length 48591.496 (873.946) 48578.821 (867.718) 56132.760 (368.832) 46623.663 (1565.744) 47630.890 (810.373) 45714.911 (1062.420) 49188.464 (753.205) 32881.580 (3110.734) 35777.096 (1949.538)
361237 Coverage 0.981 (0.0) 0.981 (0.0) 0.981 (0.0) 0.977 (0.008) 0.962 (0.0) 0.962 (0.0) 0.977 (0.008) 0.965 (0.008) 0.927 (0.031) 0.981 (0.0)
Length 47.738 (0.542) 47.440 (0.959) 30.785 (0.037) 26.208 (0.897) 30.554 (0.561) 27.182 (0.803) 35.982 (0.619) 67.660 (6.380) 18.214 (0.436) 26.897 (0.515)
361241 Coverage 0.979 (0.001) 0.978 (0.001) 0.976 (0.001) 0.978 (0.001) 0.978 (0.0) 0.964 (0.002) 0.977 (0.0) 0.972 (0.002) 0.958 (0.003) 0.979 (0.0)
Length 21.772 (0.085) 23.089 (0.106) 21.543 (0.009) 21.454 (0.109) 20.862 (0.088) 19.291 (0.060) 21.905 (0.041) 40.082 (0.045) 17.765 (0.329) 19.897 (0.062)
361242 Coverage 0.977 (0.003) 0.978 (0.001) 0.975 (0.002) 0.968 (0.003) 0.973 (0.004) 0.955 (0.002) 0.971 (0.002) 0.975 (0.0) 0.936 (0.001) 0.977 (0.001)
Length 83.892 (0.143) 99.811 (0.866) 65.672 (0.212) 68.119 (0.187) 68.155 (0.357) 80.032 (0.354) 85.678 (0.371) 53.549 (0.585) 69.388 (0.128)
361243 Coverage 0.985 (0.007) 0.985 (0.007) 0.985 (0.007) 0.956 (0.022) 0.985 (0.007) 0.97 (0.015) 0.985 (0.007) 0.978 (0.007) 0.748 (0.126) 0.985 (0.007)
Length 92.698 (1.567) 87.949 (0.147) 88.993 (2.345) 84.569 (0.557) 79.879 (0.739) 87.950 (0.308) 137.673 (15.214) 46.504 (12.957) 79.976 (2.585)
361244 Coverage 0.974 (0.015) 0.974 (0.015) 0.974 (0.015) 0.974 (0.015) 0.974 (0.015) 0.974 (0.015) 0.974 (0.015) 0.989 (0.022) 0.963 (0.037) 0.974 (0.015)
Length 5.274 (0.004) 5.274 (0.004) 5.336 (0.023) 5.284 (0.010) 5.272 (0.000) 5.269 (0.003) 5.289 (0.002) 11.283 (1.039) 0.287 (0.008) 0.287 (0.008)
361247 Coverage 0.98 (0.0) 0.976 (0.003) 0.969 (0.004) 0.974 (0.001) 0.977 (0.003) 0.945 (0.004) 0.971 (0.003) 0.982 (0.001) 0.906 (0.003) 0.974 (0.004)
Length 0.029 (0.000) 0.042 (0.000) 0.009 (0.000) 0.019 (0.000) 0.018 (0.000) 0.025 (0.000) 0.016 (0.000) 0.008 (0.000) 0.012 (0.000)
361249 Coverage 0.981 (0.003) 0.981 (0.003) 0.984 (0.002) 0.991 (0.002) 0.981 (0.003) 0.976 (0.002) 0.981 (0.003) 0.971 (0.007) 0.962 (0.011) 0.977 (0.003)
Length 3.645 (0.023) 3.674 (0.026) 3.600 (0.000) 3.322 (0.019) 3.402 (0.005) 3.201 (0.019) 3.543 (0.000) 6.533 (0.252) 2.719 (0.164) 2.972 (0.064)
Appendix H CSA Prediction Region Visualizations
We now visualize some of the prediction regions corresponding to some of the trials run in Appendix G. While we find these intervals to be connected across these tasks, we expect visualizations over multivariate output spaces, i.e. for 2D regression problems, would reveal sets to be non-connected.
Figure 9: Prediction regions across methods for task 361235 for $\alpha=0.05$ (left) and $0.025$ (right).
Figure 10: Prediction regions across methods for task 361242 for $\alpha=0.05$ (left) and $0.025$ (right).
Appendix I UCI Results
We consider those regression tasks from the UCI repository [44] that have at least 1,000 samples. The complete collection of results is presented in Table 8. As discussed in the main text, across nearly all the UCI benchmark tasks, we find that the conformalized random forest does optimally and that ensembling methods provide no further benefit over simply taking this single predictor. In this degenerate case, we would expect an optimal aggregator to simply then return this optimal single predictor. We then find CSA to consistently significantly outperform other aggregation strategies and to return a prediction region of size comparable to that of the nominal random forest.
Table 8: Average coverages across tasks for $\alpha=0.05$ are shown in the top row and average prediction set lengths in the bottom row, where both were assessed over a batch of i.i.d. test samples (10% of the dataset size). We are highlighting the robustness compared to other aggregation strategies here and so bold the best performing amongst the aggregation methods. Standard deviations and means were computed across 5 randomizations of draws of the training, calibration, and test sets. Note that, while the single-stage prediction regions are the smallest, they fail to achieve the desired coverage level and are, therefore, precluded from comparison.
Dataset Metrics Linear Model LASSO Random Forest XGBoost $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ Ensemble CSA (Single-Stage) CSA
airfoil Coverage 0.966 (0.011) 0.926 (0.011) 0.934 (0.026) 0.966 (0.011) 0.945 (0.021) 0.916 (0.016) 0.934 (0.026) 0.958 (0.032) 0.805 (0.058) 0.953 (0.011)
Length 19.062 (0.615) 11.368 (0.188) 11.770 (0.261) 12.371 (0.131) 16.227 (0.025) 16.097 (1.052) 8.609 (0.719) 14.075 (0.734)
bike Coverage 0.944 (0.007) 0.947 (0.004) 0.954 (0.002) 0.958 (0.003) 0.938 (0.006) 0.908 (0.0) 0.946 (0.003) 0.955 (0.0) 0.963 (0.007) 0.947 (0.003)
Length 3.178 (0.011) 3.343 (0.021) 0.065 (0.000) 0.111 (0.000) 0.111 (0.001) 1.722 (0.007) 0.682 (0.004) 0.156 (0.019) 0.134 (0.007)
concrete Coverage 0.977 (0.008) 0.977 (0.008) 0.981 (0.0) 0.915 (0.023) 0.962 (0.0) 0.915 (0.023) 0.981 (0.0) 0.915 (0.023) 0.854 (0.054) 0.977 (0.008)
Length 44.295 (0.424) 44.470 (0.326) 26.053 (2.114) 26.488 (1.141) 32.455 (1.165) 19.712 (6.592) 25.302 (3.244)
kin40k Coverage 0.949 (0.002) 0.949 (0.001) 0.946 (0.002) 0.948 (0.003) 0.945 (0.002) 0.919 (0.002) 0.949 (0.0) 0.945 (0.002) 0.907 (0.004) 0.941 (0.006)
Length 3.781 (0.017) 3.781 (0.016) 2.333 (0.026) 3.343 (0.026) 3.269 (0.018) 3.291 (0.009) 4.985 (0.003) 2.138 (0.001) 2.456 (0.042)
parkinsons Coverage 0.937 (0.003) 0.946 (0.0) 0.958 (0.009) 0.953 (0.007) 0.936 (0.001) 0.904 (0.008) 0.936 (0.005) 0.955 (0.004) 0.873 (0.043) 0.951 (0.003)
Length 35.957 (0.323) 36.430 (0.328) 3.254 (0.423) 11.268 (0.114) 10.992 (0.074) 21.470 (0.003) 14.161 (0.083) 3.457 (0.727) 4.584 (0.637)
pol Coverage 0.944 (0.001) 0.942 (0.002) 0.951 (0.004) 0.955 (0.001) 0.938 (0.001) 0.909 (0.003) 0.946 (0.001) 0.953 (0.005) 0.884 (0.009) 0.952 (0.003)
Length 97.944 (0.150) 97.771 (0.386) 28.000 (0.000) 48.572 (1.279) 45.056 (0.821) 69.078 (0.362) 57.432 (0.663) 24.321 (1.370) 33.230 (1.569)
protein Coverage 0.958 (0.001) 0.957 (0.002) 0.953 (0.003) 0.959 (0.003) 0.957 (0.003) 0.928 (0.003) 0.953 (0.003) 0.955 (0.0) 0.91 (0.002) 0.963 (0.004)
Length 2.316 (0.000) 2.412 (0.011) 2.151 (0.014) 2.210 (0.006) 2.134 (0.002) 2.269 (0.001) 3.707 (0.039) 1.717 (0.036) 1.994 (0.000)
pumadyn32nm Coverage 0.961 (0.011) 0.961 (0.01) 0.947 (0.001) 0.962 (0.003) 0.96 (0.007) 0.935 (0.003) 0.95 (0.013) 0.957 (0.013) 0.906 (0.026) 0.963 (0.001)
Length 3.997 (0.024) 3.979 (0.031) 1.525 (0.027) 3.518 (0.058) 3.507 (0.051) 2.350 (0.043) 3.242 (0.033) 5.351 (0.139) 1.564 (0.048) 1.858 (0.078)
tamielectric Coverage 0.953 (0.002) 0.953 (0.002) 0.947 (0.003) 0.953 (0.003) 0.952 (0.003) 0.926 (0.007) 0.949 (0.0) 0.948 (0.003) 0.909 (0.003) 0.949 (0.002)
Length 0.950 (0.001) 0.951 (0.001) 1.271 (0.006) 0.953 (0.001) 0.948 (0.001) 1.029 (0.003) 4.539 (0.038) 0.774 (0.005) 0.799 (0.004)
wine Coverage 0.96 (0.005) 0.943 (0.01) 0.931 (0.015) 0.923 (0.02) 0.928 (0.005) 0.884 (0.015) 0.948 (0.005) 0.946 (0.015) 0.805 (0.054) 0.933 (0.015)
Length 2.352 (0.002) 3.521 (0.025) 2.619 (0.050) 2.621 (0.039) 3.390 (0.042) 1.683 (0.224) 2.291 (0.026)
Appendix J Robust Traffic Routing Setup
@@ Appendix I UCI Results
Appendix K Compute Details
All OpenML were all run on a standard-grade CPU. The deep learning-based experiments, namely the ImageNet classification and traffic forecasting predict-then-optimize task, were performed on an Nvidia RTX 2080 Ti GPU. Such experiments, however, were conducted with publicly available, pre-trained models provided by the works respectively referenced in the sections describing the experimental setups.
Appendix L Conformal Aggregation Methods
We now describe the methods from [25] that were compared against experimentally, specifically the standard majority-vote $\mathcal{C}^{M}$, partially randomized thresholding $\mathcal{C}^{R}$, and fully randomized thresholding $\mathcal{C}^{U}$ approaches. As discussed in Section 2.4, these methods all follow the structural form of
and largely differ in their choice of weights and thresholds. The standard majority-vote $\mathcal{C}^{M}$ is the most natural choice, defined by